# MedPlan: A Two-Stage RAG-Based System for Personalized Medical Plan Generation

### v1

@@ Abstract
Despite recent success in applying large language models (LLMs) to electronic health records (EHR), most systems focus primarily on assessment rather than treatment planning. We identify three critical limitations in current approaches: they generate treatment plans in a single pass rather than following the sequential reasoning process used by clinicians; they rarely incorporate patient-specific historical context; and they fail to effectively distinguish between subjective and objective clinical information. Motivated by the SOAP methodology (Subjective, Objective, Assessment, Plan), we introduce MedPlan, a novel framework that structures LLM reasoning to align with real-life clinician workflows. Our approach employs a two-stage architecture that first generates a clinical assessment based on patient symptoms and objective data, then formulates a structured treatment plan informed by this assessment and enriched with patient-specific information through retrieval-augmented generation. Comprehensive evaluation demonstrates that our method significantly outperforms baseline approaches in both assessment accuracy and treatment plan quality.
@@
Hsin-Ling Hsu1thanks: Equal contribution, Cong-Tinh Dao211footnotemark: 1, Luning Wang3, Zitao Shuai3, Thao Nguyen Minh Phan2, Jun-En Ding 4, Chun-Chieh Liao4, Pengfei Hu 4, Xiaoxue Han 4, Chih-Ho Hsu 5, Dongsheng Luo 6, Wen-Chih Peng 2, Feng Liu 4, Fang-Ming Hung 5, Chenwei Wu 3 1National Chengchi University, 2National Yang Ming Chiao Tung University, 3University of Michigan, 4Stevens Institute of Technology, 5Far Eastern Memorial Hospital 6Florida International University Correspondence: chenweiwu99@gmail.com
@@ 1 Introduction
Deploying large language models (LLMs) for electronic health records (EHR) Evans (2016) analysis in high-stakes medical environments presents significant opportunities for enhancing patient care through automation and improved clinical decision support Yang et al. (2022); Zhang et al. (2024); Sakai and Lam (2025). Despite recent progress in adapting LLM to medical domain Tang et al. (2025); Jiang et al. (2025), most existing LLM systems Palepu et al. (2025) for EHR focus largely on diagnostic assessment tasks, neglecting the crucial subsequent step of structured, patient-specific treatment planning Sarker et al. (2021); Curtis et al. (2017). Effective LLM-based planning could significantly reduce physician cognitive load, standardize care protocols, decrease treatment variability, and enable more personalized interventions.
@@
To address these challenges, we introduced MedPlan, a novel framework that explicitly structures LLM reasoning to mirror the SOAP clinical workflow. Our approach operates in two clinically-grounded stages that parallel physician cognitive processes: (1) a diagnostic phase where we generate an assessment (A) based on patient symptoms and clinical data (S and O), completing the diagnostic reasoning before proceeding, and (2) a therapeutic phase where we formulate a structured treatment plan (P) directly informed by the assessment and tailored to patient-specific factors. This two-stage architecture faithfully replicates how clinicians reason—first establishing what is happening before determining what should be done. We enhanced the planning phase through patient-specific retrieval-augmented generation (RAG), allowing the model to consider longitudinal patient information—mirroring how physicians integrate medical history into their treatment decisions.
@@ 3 Methodology
To obtain accurate and personalized clinical plannings that align with physician workflow, we present MedPlan, a trustworthy clinical decision support system that employs a two-stage generation pipeline, mirroring the natural progression of clinical planning. To get high-quality planning, we propose to first generate an assessment based on the patient data, then create the treatment plan based on both the patient data and the generated assessment. This separation follows the established SOAP protocol, where clinicians first analyze symptoms and findings to form a diagnosis before determining appropriate interventions. We also explicitly separate S and O components in our prompts (see Appendix D), allowing the model to distinctly process patient-reported symptoms versus clinical observations—a key distinction that enhances clinical relevance. To enhance the personalization and accuracy of the generated plannings, we further leverage two types of references during generation: (1) self-history references—the patient’s previous SOAP records, and (2) cross-patient references—similar cases from other patients. Specifically, for the $i$-th patient, we retrieve their latest $N_{\text{hist}}$ SOAP records as self-history references, formulated as $\mathcal{R}^{\text{hist}}{i}={(S{j},O_{j},A_{j},P_{j})\mid j\in{1,2,...,N_{\text{hist}}}}$. Furthermore, to better align with the clinical reasoning patterns, we incorporate instruction tuning on the models that generate A and P before deploying our two-stage pipeline. Figure 2 illustrates the overall architecture of our inference workflow.
@@ 3.1 Assessment Generation Stage
In the Assessment Generation Stage, we integrated the patient’s current $S$ and $O$ information with both self-history references $\mathcal{R}_{\text{hist}}$ and cross-patient references $\mathcal{R}^{SOA}=\{(S_{j},O_{j},A_{j})\}_{j=1}^{N_{\text{ref}}}$. To identify the most relevant cross-patient references, we employ a two-step retrieval process. First, we retrieve $N_{\text{sim}}$ candidate references $\mathcal{R}^{\text{SOA}}_{\text{sim}}$ via semantic similarity search using a bi-encoder model Karpukhin et al. (2020), which efficiently computes embeddings for the query and candidates separately. Then, we refined this selection using a more computationally intensive but more accurate cross-encoder re-ranking model Nogueira and Cho (2020) that evaluates the fine-grained clinical relevance by jointly encoding the query and each candidate:
@@ 3.2 Plan Generation Stage
In the Plan Generation Stage, we utilized the generated assessment $A_{\text{gen}}$ along with the original $S$ and $O$ to retrieve and generate an appropriate treatment plan. Mirroring the clinical practice where physicians formulate treatment plans based on their diagnostic assessment and patient information, we employed another retrieval process to find relevant plan references $\mathcal{R}^{SOAP}=\{(S_{j},O_{j},A_{j},P_{j})\}_{j=1}^{N_{\text{ref}}}$. Similar to the previous stage, we use a two-step retrieval approach. First, we retrieved $N_{\text{sim}}$ candidate references $\mathcal{R}^{\text{SOAP}}_{\text{sim}}$ via semantic similarity search using a bi-encoder model. Then, we refined this selection using a cross-encoder re-ranking model:
@@ 4.1 Datasets
This study utilized 350,684 outpatient and emergency EHR SOAP notes from 55,890 patients collected at a Eastern Memorial Hospital (FEMH) in 2021. All data were de-identified prior to analysis. We preprocessed all SOAP notes by removing records shorter than two characters and normalizing text (eliminating newlines, redundant spaces, and consecutive punctuation).
@@ 4.2 Metrics
Table 1: Performance comparison of different models and settings
@@ 4.3 Implementation Details
We utilized prompt engineering techniques and employed LoRA for parameter-efficient fine-tuning. We determined that $N_{\text{hist}}=20$ and $N_{\text{ref}}=10$ for RAG, with $N_{\text{sim}}=80$ initial candidates for semantic similarity search. To evaluate MedPlan, we simulated clinical diagnostic processes by using the first $N-2$ visits as history $\mathcal{R}_{\text{hist}}$ and the second-to-last visit as a training target for patients with $N$ visits. During testing, we used the first $N-1$ visits as history $\mathcal{R}_{\text{hist}}$ with the most recent visit as the evaluation target. We experimented with various configurations by enabling different components in our pipeline, including: Self-history, Instruction Tuning, Cross-patient References, Direct Plan Generation, and a Two-step Approach with Pre-plan Assessment. Each experiment involved selecting a subset of these components. Additional implementation details are in Appendix B.1.
@@ 5 Clinical Application Demo and System Design
To demonstrate the real-world applicability of our Plan generation system, we developed a clinical prototype that has been reviewed by practicing physicians for viability in actual healthcare settings. Our system works as follows: The physician first inputs the patient’s S and O, and the system generates A and P based on these inputs. At the same time, physicians can modify A according to their clinical judgment and regenerate P, while our system can update retrievals through RAG, which leverages a knowledge base of patient SOAP notes. The more specific technical architecture of the backend system is shown in Figure 2. The frontend is developed using React, the backend is based on FastAPI service, and communication between frontend and backend is conducted through RESTful API. The core of the system includes two specialized LLMs, responsible for generating A and P respectively. The system uses Microsoft SQL (MSSQL) database to store patient historical data, and enhances semantic retrieval and case matching through vector embedding using Weaviate database.
More detailed framework can be found in Appendix B.
@@ 6 Conclusion
Appendix A System Demonstration
The personal medical plan generation system will be demonstrated online soon. The demo site is hosted under a domain with full WHOIS privacy protection and contains no identifying information about the authors.
Appendix B System Architecture
@@
B.1 Implementation Details
We instruction-tuned multiple LLMs, including Medical-Llama3-8B, Medical-Mixtral-7B-v2k, and Bio-Medical-Llama3-8B, using the Unsloth Daniel Han and team (2023) framework for optimized training. It is performed on NVIDIA RTX 6000 Ada Generation, with LoRA (Low-Rank Adaptation) dynamically allocated based on model selection. The models are trained with a sequence length of 65,536 tokens, allowing long-context processing. We employed the AdamW optimizer with 8-bit precision, using a cosine learning rate scheduler with warmup steps (0.016 $\times$ total training steps).
Appendix C Generation Samples
@@
When a subset of these generation samples was presented to physicians at a FEMH for evaluation, the proposed method demonstrated approximately 66% better performance in clinical assessments compared to the baseline approach.
@@
Appendix D Prompt Template

### v2

@@ Abstract
Despite recent success in applying large language models (LLMs) to electronic health records (EHR), most systems focus primarily on assessment rather than treatment planning. We identify three critical limitations in current approaches: they generate treatment plans in a single pass rather than following the sequential reasoning process used by clinicians; they rarely incorporate patient-specific historical context; and they fail to effectively distinguish between subjective and objective clinical information. Motivated by the SOAP methodology (Subjective, Objective, Assessment, Plan), we introduce MedPlan, a novel framework that structures LLM reasoning to align with real-life clinician workflows. Our approach employs a two-stage architecture that first generates a clinical assessment based on patient symptoms and objective data, then formulates a structured treatment plan informed by this assessment and enriched with patient-specific information through retrieval-augmented generation. Comprehensive evaluation demonstrates that our method significantly outperforms baseline approaches in both assessment accuracy and treatment plan quality. Our demo system and code are available at https://github.com/JustinHsu1019/MedPlan.
@@
Hsin-Ling Hsu1thanks: Equal contribution, Cong-Tinh Dao2,311footnotemark: 1, Luning Wang4, Zitao Shuai4, Nguyen Minh Thao Phan2,3, Jun-En Ding 5, Chun-Chieh Liao5, Pengfei Hu 5, Xiaoxue Han 5, Chih-Ho Hsu 6, Dongsheng Luo 7, Wen-Chih Peng 2, Feng Liu 5, Fang-Ming Hung 6, Chenwei Wu 4 1National Chengchi University, 2National Yang Ming Chiao Tung University, 3Can Tho University, 4University of Michigan, 5Stevens Institute of Technology, 6Far Eastern Memorial Hospital 7Florida International University Correspondence: chenweiwu99@gmail.com
@@ 1 Introduction
Deploying large language models (LLMs) for electronic health records (EHR) Evans (2016) analysis in high-stakes medical environments presents significant opportunities for enhancing patient care through automation and improved clinical decision support Yang et al. (2022); Zhang et al. (2024); Sakai and Lam (2025); Ding et al. (2024). Despite recent progress in adapting LLM to medical domain Tang et al. (2025); Jiang et al. (2025); Restrepo et al. (2025), most existing LLM systems Palepu et al. (2025); Fan and Tao (2024) for EHR focus largely on diagnostic assessment tasks, neglecting the crucial subsequent step of structured, patient-specific treatment planning Sarker et al. (2021); Curtis et al. (2017). Effective LLM-based planning could significantly reduce physician cognitive load, standardize care protocols, decrease treatment variability, and enable more personalized interventions.
@@
To address these challenges, we introduced MedPlan, a novel framework that explicitly structures LLM reasoning to mirror the SOAP clinical workflow. Our approach operates in two clinically-grounded stages that parallel physician cognitive processes: (1) a diagnostic phase where we generate an assessment (A) based on patient symptoms and clinical data (S and O), completing the diagnostic reasoning before proceeding, and (2) a therapeutic phase where we formulate a structured treatment plan (P) directly informed by the assessment and tailored to patient-specific factors. This two-stage architecture faithfully replicates how clinicians reason—first establishing what is happening before determining what should be done. We enhanced the planning phase through patient-specific retrieval-augmented generation (RAG) Lewis et al. (2020), allowing the model to consider longitudinal patient information—mirroring how physicians integrate medical history into their treatment decisions.
@@ 3 Methodology
To obtain accurate and personalized clinical plannings that align with physician workflow, we present MedPlan, a trustworthy clinical decision support system that employs a two-stage generation pipeline, mirroring the natural progression of clinical planning. To get high-quality planning, we propose to first generate an assessment based on the patient data, then create the treatment plan based on both the patient data and the generated assessment. This separation follows the established SOAP protocol, where clinicians first analyze symptoms and findings to form a diagnosis before determining appropriate interventions. We also explicitly separate S and O components in our prompts (see Appendix C), allowing the model to distinctly process patient-reported symptoms versus clinical observations—a key distinction that enhances clinical relevance. To enhance the personalization and accuracy of the generated plannings, we further leverage two types of references during generation: (1) self-history references—the patient’s previous SOAP records, and (2) cross-patient references—similar cases from other patients. Specifically, for the $i$-th patient, we retrieve their latest $N_{\text{hist}}$ SOAP records as self-history references, formulated as $\mathcal{R}^{\text{hist}}{i}={(S{j},O_{j},A_{j},P_{j})\mid j\in{1,2,...,N_{\text{hist}}}}$. Furthermore, to better align with the clinical reasoning patterns, we incorporate instruction tuning on the models that generate A and P before deploying our two-stage pipeline. Figure 2 illustrates the overall architecture of our inference workflow.
@@ 3.1 Assessment Generation Stage
In the Assessment Generation Stage, we integrated the patient’s current $S$ and $O$ information with both self-history references $\mathcal{R}_{\text{hist}}$ and cross-patient references $\mathcal{R}^{SOA}=\{(S_{j},O_{j},A_{j})\}_{j=1}^{N_{\text{ref}}}$. To identify the most relevant cross-patient references, we employ a two-step retrieval process. First, we retrieve $N_{\text{sim}}$ candidate references $\mathcal{R}^{\text{SOA}}_{\text{sim}}$ via hybrid retrieval Ma et al. (2020); Bruch et al. (2023); Hsu and Tzeng (2025) combining BM25 Robertson et al. (1995) and bi-encoder semantic search Karpukhin et al. (2020), leveraging both keyword matching and semantic similarity. Then, we refined this selection using a more computationally intensive but more accurate cross-encoder re-ranking model Nogueira and Cho (2020) that evaluates the fine-grained clinical relevance by jointly encoding the query and each candidate:
@@ 3.2 Plan Generation Stage
In the Plan Generation Stage, we utilized the generated assessment $A_{\text{gen}}$ along with the original $S$ and $O$ to retrieve and generate an appropriate treatment plan. Mirroring the clinical practice where physicians formulate treatment plans based on their diagnostic assessment and patient information, we employed another retrieval process to find relevant plan references $\mathcal{R}^{SOAP}=\{(S_{j},O_{j},A_{j},P_{j})\}_{j=1}^{N_{\text{ref}}}$. Similar to the previous stage, we use a two-step retrieval approach. First, we retrieve $N_{\text{sim}}$ candidate references $\mathcal{R}^{\text{SOAP}}_{\text{sim}}$ via hybrid retrieval combining BM25 and bi-encoder semantic search. Then, we refined this selection using a cross-encoder re-ranking model:
@@ 4.1 Datasets
This study utilized 350,684 outpatient and emergency EHR SOAP notes from 55,890 patients collected at Far Eastern Memorial Hospital (FEMH) in 2021. All data were de-identified prior to analysis. We preprocessed all SOAP notes by removing records shorter than two characters and normalizing text (eliminating newlines, redundant spaces, and consecutive punctuation).
@@ 4.2 Metrics
Table 1: Performance Comparison of Different Models and Settings for Plan Generation
@@ 4.3 Implementation Details
We utilized prompt engineering techniques and applied LoRA for parameter-efficient fine-tuning. Specifically, we instruction-tuned several open-source LLMs—Medical-Llama3-8B Vsevolodovna (2024a), Medical-Mixtral-7B-v2k Vsevolodovna (2024b), and Bio-Medical-Llama3-8B ContactDoctor (2024)—using the Unsloth framework Daniel Han and team (2023). To support long-context retrieval in our RAG-based design, we adopted OpenAI’s text-embedding-3-large model OpenAI (2024) for semantic similarity search, and used VoyageAI Reranker-2 VoyageAI (2024) as a cross-encoder model to re-rank the retrieved candidates. For baseline comparison, we additionally evaluated two general-purpose models: o1 OpenAI (2024b) and GPT-4o OpenAI (2024a), without domain-specific adaptation.
We set $N_{\text{hist}}=20$ and $N_{\text{ref}}=10$ for our RAG module, retrieving $N_{\text{sim}}=80$ initial candidates based on semantic similarity. To evaluate MedPlan, we simulated clinical diagnostic processes by using the first $N{-}2$ visits as history $\mathcal{R}_{\text{hist}}$ and the second-to-last visit as the training target for patients with $N$ visits, while the first $N{-}1$ visits and the most recent visit were used as history and evaluation target respectively during testing. We conducted ablation experiments with various configurations by selectively enabling components in our pipeline, including: Self-history, Instruction Tuning, Cross-patient References, Direct Plan Generation, and a Two-step Approach with Pre-plan Assessment. Additional implementation details, including training environment and hyperparameter settings, are provided in Appendix A.1.
@@ 5 Clinical Application Demo and System Design
To demonstrate the real-world applicability of our Plan generation system, we developed a clinical prototype that has been reviewed by practicing physicians for viability in actual healthcare settings. An overview of the clinical interface is shown in Figure 4. Our system works as follows: The physician first inputs the patient’s S and O, and the system generates A and P based on these inputs. At the same time, physicians can modify A according to their clinical judgment and regenerate P, while our system can update retrievals through RAG, which leverages a knowledge base of patient SOAP notes. The more specific technical architecture of the backend system is shown in Figure 2. The frontend is developed using React, the backend is based on FastAPI service, and communication between frontend and backend is conducted through RESTful API. The core of the system includes two specialized LLMs, responsible for generating A and P respectively. The system uses Microsoft SQL (MSSQL) database to store patient historical data, and enhances semantic retrieval and case matching through vector embedding using Weaviate database.
The detailed system architecture is provided in Appendix A.
@@ 6 Conclusion
Appendix A System Architecture
@@
A.1 Implementation Details
We instruction-tuned three domain-specific LLMs—Medical-Llama3-8B Vsevolodovna (2024a), Medical-Mixtral-7B-v2k Vsevolodovna (2024b), and Bio-Medical-Llama3-8B ContactDoctor (2024)—using the Unsloth framework Daniel Han and team (2023) for efficient adaptation with long-context support. All models were trained on NVIDIA RTX 6000 Ada Generation GPUs with Low-Rank Adaptation (LoRA), dynamically adjusted for each model’s architecture. A maximum sequence length of 65,536 tokens was used to accommodate extended patient histories and cross-patient references. The training employed the AdamW optimizer in 8-bit precision, along with a cosine learning rate scheduler and a warm-up phase equal to 1.6% of the total steps.
For semantic retrieval, we used OpenAI’s text-embedding-3-large model OpenAI (2024), which supports high-dimensional dense representations suitable for medical content. As our cross-encoder model, we employed the VoyageAI Reranker-2 VoyageAI (2024), which was used to re-rank the semantically retrieved candidates in our RAG pipeline. All experiments were conducted under consistent hardware and software configurations to ensure comparability.
Appendix B Generation Samples
@@
When a subset of the generated samples was presented to physicians at Far Eastern Memorial Hospital (FEMH) for evaluation, the proposed method demonstrated approximately 66% improvement in clinical assessments compared to the baseline approach.
@@
Appendix C Prompt Template